OpenStudy (anonymous):

solve the inequality. x-5 < -3

OpenStudy (anonymous):

can you solve this: x-5=-3 ? :)

OpenStudy (anonymous):

Basically get the x variable alone..to do that, add 5 to both sides...making it x<2.

OpenStudy (anonymous):

If it needs two solutions, break ti up into two equations...x-5 < -3 and x-5>3. Then solve for x in each.

OpenStudy (anonymous):

i got 2 x-5<-3 +5 +5 x<2

OpenStudy (anonymous):

but he doesn't have x-5 in absolute value.. that's not necessary ...

OpenStudy (anonymous):

thank you mads4566

OpenStudy (anonymous):

The problem to solve is: x-5 ‹ -3 First, let's work on the left hand side of your inequality, the x-5 This means, for instance, to see if it can be simplified at all. x-5 evaluates to x-5 So, all-in-all, the left hand side of your inequality can be written as: x-5 Now, let's work on the right hand side of your inequality, the -3 Because of the minus sign 3 becomes - 3 The answer is -3 The right hand side of your inequality can be written as: -3 So with these (any) simplifications, the inequality we'll set out to solve is: x-5 ‹ -3 Move the -5 to the right hand side by adding 5 to both sides, like this: To the left hand side: -5 + 5 = 0 The answer is x To the right hand side: -3 + 5 = 2 The answer is 2 Now, the inequality reads: x ‹ 2 To isolate the x, we have to divide both sides of the inequality by the other "stuff" (variables or coefficients) around the x on the left side of the inequality. and this is the final solution to your inequality.